### v3

Abstract
Tidal energy is one of the key components in increasing the penetration of renewable energy. High tidal energy penetration into the electrical grid depends on accurate tidal current speed forecasting. Model inaccuracies hinder forecast accuracy. Previous research primarily used physical models to forecast tidal current speed, yet tidal current variations influenced by the orbital periods of celestial bodies make accurate physical modeling challenging. Research on the multi-periodicity of tides is crucial for forecasting tidal current speed. We propose the Wavelet-Enhanced Convolutional Network to learn multi-periodicity. The framework embeds intra-period and inter-period variations of one-dimensional tidal current data into the rows and columns, respectively, of a two-dimensional tensor. Then, the two-dimensional variations of the sequence can be processed by convolutional kernels. We integrate a time-frequency analysis method into the framework to further address local periodic features. Additionally, to enhance the framework’s stability, we optimize the framework’s hyperparameters with the Tree-structured Parzen Estimator. The proposed framework captures multi-periodic dependencies in tidal current data. Numerical results show a 10-step average Mean Absolute Error of 0.025, with at least a 1.18% error reduction compared to other baselines. Further ablation studies show a 1.4% reduction in Mean Absolute Percentage Error on the data with artificially added periodic fluctuations.
keywords:
[…]
1 Introduction
1.1 Background
The development of renewable energy can promote emission reduction [1]. Tidal energy is considered one of the most competitive forms of renewable energy [2].
Tidal energy attracts the attention of many coastal countries [3], especially the UK. The potential of tidal energy lies in three aspects: cost-effectiveness, stability, and predictability. As tidal energy gains more attention, its cost-effectiveness is expected to improve, as shown in Figure 1. Tidal energy may replace other renewable energies when the Levelized Cost of Energy (LCOE) falls below £49-55/MWh (expected by 2050) [4]. Additionally, tidal energy offers high stability, which is crucial for the economic feasibility of any renewable energy project [5]. Therefore, the introduction of tidal energy can enhance grid stability. Moreover, tidal energy provides the advantage of predictability over large time scales. Consequently, tidal energy is an excellent choice as a base-load energy supply [2].
Tidal current turbines generate electricity using the kinetic energy of tidal currents, like wind turbines use the kinetic energy of air [6]. Therefore, fluctuations in Tidal Current Speed (TCS) reduce grid stability [7]. Enhancing the accuracy of short-term forecasting helps mitigate the impact of TCS fluctuations on grid stability, thereby increasing the penetration of tidal energy into the grid [8]. Research [9] proposes to explore the statistical properties of real-life time series and design customized deep learning models for high-performance time series forecasting. The research on statistical properties of time series inspires us to explore the multi-periodicity within the TCS series. We increase the consideration of the statistical properties of TCS to improve the accuracy of TCSF.
Figure 1: Evaluation of Levelized Cost of Energy (LCOE) and Cumulative Installed Capacity in the UK. ±1 SD indicates data variation within one standard deviation around its expected value.
Nomenclature
$\beta$ Output length $\boldsymbol{\omega}$ A hyperparameter combination $\boldsymbol{\omega}_{\text{space}}$ Hyperparameter combination search space $\chi(\cdot)$ Probability density function for hyperparameter configurations where $y$ is better than the threshold $y^{*}$ $\gamma(\cdot)$ Probability density function for hyperparameter configurations where $y$ is worse than the threshold $y^{*}$ $\hat{\mathbf{X}}^{l,i}_{\text{2D}}$ 2D tensor processed by Inception $\hat{\mathbf{X}}^{l,i}_{\text{1D}}$ 1D sequence after reshaping back $\hat{\textbf{{x}}}^{n}(t)$ The forecasted value $\hat{A}^{l-1}_{f_{1}}$ Weights for sequence aggregation $\textbf{{cD}}_{j}$ Average detail coefficients at the $j$-th level for all sequences $\textbf{{x}}^{n}(t)$ The time point on the $n$-th sequence in $\textbf{{X}}_{\text{1D}}$ $\textbf{{X}}_{\text{1D}}$ Multivariate input time series $\textbf{{X}}_{\text{1D}}^{l}$ Features of the $l$-th layer of $X_{\text{1D}}$ $\textbf{{X}}_{\text{2D}}^{l,i}$ 2D tensor obtained by reshaping $\xi_{j}$ Average amplitude at the $j$-th level $A_{f_{i}}$ Amplitude corresponding to $f_{i}$ $f_{i}$ The $i$-th frequency $k$ Top $k$ amplitudes are selected NN Number of sequences $p_{i}$ The $i$-th period $T$ Number of sampling points per sequence $y$ Loss function BSL Below Sea Level DWT Discrete Wavelet Transform FFT Fast Fourier Transform MAE Mean Absolute Error MAPE Mean Absolute Percentage Error MSE Mean Squared Error MSPE Mean Squared Percentage Error RS Random Sampler TCS Tidal Current Speed TCSF Tidal Current Speed Forecasting TPE Tree-structured Parzen Estimator WECN Wavelet-Enhanced Convolutional Network
1.2 Literature review
Previous research focused on employing various models for TCSF. Existing tidal current forecasting models are divided into five categories: tidal harmonic models, physical models, statistical models, traditional machine learning models, and deep learning models. The summary of the advantages and disadvantages of the mentioned models are listed in Table 1.
Forecasting model
Without Dependence on Specific Locations
Considering local variations
Without Complex Feature Engineering
Considering Multi-Periodicity
Table 1: The summary of the advantages and disadvantages of the mentioned models. The symbols ①-⑥ represent Tidal harmonic models, Physical models, Statistical models, Traditional machine learning models, Deep learning models, and Our models. respectively. Additionally, indicates the model has the corresponding feature, while indicates it does not.
Tidal harmonic models
In 1883, Sir G. H. Darwin proposed the harmonic analysis method [10]. He stated, “The tidal oscillations of the ocean can be represented as the sum of several simple harmonic motions.” The harmonic analysis method was later developed by Doodson, who established the famous tidal harmonic component theory. Harmonic models rely on parameter estimation. Doodson continued this research and proposed estimating the parameters of the harmonic model with the least squares method [11]. Earlier technologies did not allow equipment to measure TCS. Research [12] applied harmonic models for TCSF latter. The harmonic model takes into account the multi-periodicity of TCS. However, the harmonic model cannot adapt to the local periodic changes of TCS. There were more models for TCSF as computing power increased.
Physical models
The emergence of supercomputers brought advancements to physical models. Based on differential equations, physical models use two-dimensional or three-dimensional hydrodynamic models to make forecasts [13]. Physical models can adapt to the non-stationarity of TCS. However, modeling with differential equations depends on physical knowledge. Additionally, solving these equations requires continuous measurement of boundary conditions. Therefore, data-driven models are needed for TCSF.
Statistical models
Statistical models are suitable as a data-driven approach for TCSF. They avoid modeling based on specific TCS measurement locations. However, statistical models rely on strict stationarity assumptions. Therefore, statistical methods have some limitations and they have relatively high requirements in terms of data size, stability of time series, data distribution, and so on [14].
Machine learning models
With the rise of artificial intelligence technology, a large number of forecasting methods based on machine learning technology have emerged [15]. Research [16, 17, 18] indicates that machine learning models outperform both statistical and harmonic models in predictive performance. However, machine learning models cannot determine data features. Feature extraction usually requires intuitive knowledge for quantification and expansion.
Deep learning models
Deep learning models can provide accurate renewable energy generation forecasts [19]. Deep learning models are applied in the renewable energy field to reduce the complexity of feature engineering. Research [20] used Long Short-Term Memory (LSTM) neural networks for tidal level forecasting. However, deep learning models are still underutilized in the tidal energy field. In other words, while deep learning methods have certain advantages, they are not designed for tidal current time series.
1.3 Research gaps and proposed solution
Existing research still has limitations. Tidal current time series demonstrate multi-periodicity [21]. These time series include multiple periodic components, covering both astronomical cycles and local periods. Astronomical cycles include diurnal and semi-diurnal tides, among others. In addition to astronomical cycles, there are also local periodicities in tidal currents, manifested as local variations. For example, interactions between tides and waves cause local changes [22]. Multiple overlapping and interacting cycles complicate the model of tidal current time series. Multi-periodicity presents an obstacle to TCSF. Secondly, existing research lacks deep learning models designed for tidal current time series. Therefore, using deep learning methods to consider the multi-periodicity in tidal current time series has sparked our research interest.
We are inspired by the TimesNet model proposed by Wu et al. [23]. The model focuses on multi-periodicity. The model uses the Fast Fourier Transform (FFT) to extract multiple periods from the time series. Based on the periods, the model transforms each 1D sequence into a 2D tensor. Then, the model uses visual backbone networks to capture the 2D variations in the time series. However, the framework lacks consideration for the local periodicity of tidal currents. Research [24] used time series decomposition methods to handle time series. The research inspired us: Discrete Wavelet Transform (DWT) can provide local information on signals in both time and frequency domains. We can handle the series at different scales to capture local periodicity. Additionally, the DWT module helps address local periodicity. The key challenge lies in integrating local periodicity into the learning of multi-periodicity. We propose a framework, Wavelet-Enhanced Convolutional Network (WECN), to address the aforementioned challenges. We use DWT to decompose the time series and measure sub-band amplitudes to identify dominant periods, feeding them into the predictor without reconstructing components. Finally, we consider the forecasting process to be an optimization problem. However, as the dimensionality of hyperparameters increases, the Gaussian model becomes increasingly complex. The Tree-structured Parzen Estimator (TPE) algorithm optimizes hyperparameters to handle the high dimensionality of the hyperparameter space [25].
1.4 Novelty and contributions
Overall, the contributions of this paper are as follows:
  • We designed a time series forecasting framework for learning multi-periodicity. The framework outperforms existing models in terms of metrics.
  • We use DWT to account for local periodicity, aiming for better performance in multi-periodicity learning. Experiments validate the framework’s performance.
  • We use TPE for hyperparameter optimization to improve forecasting accuracy. Ablation experiments demonstrate the effectiveness of the method.
  • We design a novel framework for forecasting tidal current speed. The framework can contribute to advancements in the field of tidal energy.
1.5 Paper organization
This research is organized as follows: Section 2 shows the description of the multi-periodicity and the proposed framework. Section 3 presents case studies on the tidal data. Section 4 gives the conclusions and future works.
2 Methodology
We construct a framework that considers the multi-periodicity of TCS. Section 2.1 shows the process of extracting periods using the improved wavelet. The periods extracted by DWT are used for the predictor. Section 2.2 shows the predictor used in this research. Section 2.3 shows the process of hyperparameter optimization.
Without loss of generality, a multivariate time series $\textbf{{X}}_{\text{1D}}=[\textit{{x}}^{1},\textit{{x}}^{2},\cdots,\textit{{x}}^{n},\cdots,\textit{{x}}^{N}]$ is defined, where $\textbf{{X}}_{\text{1D}}$ is a two-dimensional array of shape $(T,N)$. The shape of $(T,N)$ represents $N$ time series, each with $T$ sampling points. $\textbf{{x}}^{n}$ denotes the $n$-th time series, $\textbf{{x}}^{n}=\{\textit{x}^{n}(1),\textit{x}^{n}(2),\ldots,\textit{x}^{n}(T)\}$.
2.1 Multi-scale time series analysis
Multi-periodicity fundamentally arises from different inherent properties within the time series. The local periodicity of TCS is an intrinsic property in tidal current time series. We can analyze time series at multiple scales. We can identify information in both time and frequency domains. The approach is used to identify local periodicity. Therefore, we use the improved wavelet to extract periods for multi-scale time series analysis. We present the process of extracting periods. In this work, Discrete Wavelet Transform (DWT) is used solely for decomposition and amplitude-based measurement to select dominant scales; no inverse transform or signal reconstruction is performed.
DWT is first performed on the time series $\textbf{{x}}^{n}$. The DWT uses a set of high-pass filters $h$ and low-pass filters $g$ to decompose the time series signal into different resolutions. As the decomposition levels increase, the frequency resolution of the original signal increases. At the $j$-th level, the outputs of the high-pass and low-pass filters are represented as detail coefficients $\textbf{{cD}}^{n}_{j}$ and approximation coefficients $\textbf{{cA}}^{n}_{j}$, respectively. The detail coefficients describe short-term changes in the series, while the approximation coefficients describe the long-term trends of the signal. The initial approximation is $\textbf{{cA}}^{n}_{0}=\textbf{{x}}^{n}$. To obtain the $\textbf{{cD}}^{n}_{j}$ and $\textbf{{cA}}^{n}_{j}$ , the DWT operation at the $j$-th level is represented as:
$(\textbf{{cD}}^{n}_{j},\textbf{{cA}}^{n}_{j})=\text{DWT}(\textbf{{cA}}^{n}_{j-1})=(h*\textbf{{cA}}^{n}_{j-1},g*\textbf{{cA}}^{n}_{j-1}),$ (1)
where j=1,2,,Jj=1,2,...,J, $\text{DWT}(\cdot)$ denotes the execution of a single discrete wavelet transform, * represents the convolution operation, and the selection of $h$ and $g$ depends on the wavelet basis $w$, a hyperparameter that needs optimization.
After the time series $\textbf{{x}}^{n}$ is decomposed into $J$ levels, $J$ detail coefficients $\textbf{{cD}}^{n}_{1},\textbf{{ cD}}^{n}_{2},...,\textbf{{cD}}^{n}_{J}]$ are output. At the $j$-th level, the detail coefficients $\textbf{{cD}}^{n}_{j}$ has a length of $\frac{T}{2^{j}}$. For $\textbf{{X}}_{\text{1D}}$, each series is decomposed into the same number of levels. The information from multiple series is first merged to obtain the average detail coefficients $\textbf{{cD}}_{j}$, then the amplitude $\xi_{j}$ is calculated. $\xi_{j}$ can represent the intensity of signal change at that level, expressed as:
$\textbf{{cD}}_{j}=\frac{1}{N}\sum_{n=1}^{N}\textbf{{cD}}^{n}_{j},\quad\xi_{j}=\|\textbf{{cD}}_{j}\|^{2},$ (2)
where $j=1,2,\ldots,J$.
The detail layers with the largest amplitudes show the most significant frequency changes in the signal. Inspired by other research, to avoid meaningless noise from high frequencies, the top $k$ detail layers with the largest amplitudes are selected [26]. $k$ is one of the hyperparameters to be optimized. If $J$ is less than $k$, all amplitudes are selected to form the amplitude array $\boldsymbol{\xi}=[\xi_{1},\xi_{2},\ldots,\xi_{J}]$, and $k$ is updated to $J$. If $J$ is greater than or equal to $k$, the top $k$ values are selected to form the amplitude array $\boldsymbol{\xi}=\text{top}(\xi_{1},\xi_{2},\ldots,\xi_{J})$. Here, $\text{top}(\cdot)$ denotes the top $k$ elements selection to form a new array. Next, the index array $\psi=\text{index}(\xi)$ is obtained, where $\text{index}(\cdot)$ denotes the indices of the elements in the set. Finally, the frequency $f_{i}$ and period $p_{i}$ for the corresponding levels are calculated, expressed as:
$f_{i}=\frac{F}{2^{j}},\quad p_{i}=\left[\frac{1}{f_{i}}\right],$ (3)
where $i=1,2,\ldots,k$, $F$ denotes the sampling frequency of the time series data and $j\in\psi$.
Equation (3) results in the frequency $f_{i}$ and corresponding period $p_{i}$. To associate the amplitudes with their frequencies, the amplitudes are re-expressed as $\boldsymbol{A}=[A_{f_{1}},\ldots,A_{f_{k}}]$.
Figure 2: Flowchart of the Period Extraction Module, illustrating the process of transforming multivariate time series inputs into the main period ($P$) and frequency ($f$).
2.2 Tidal current speed predictor
We use WECN to forecast after we obtain the periodic information. Existing methods have limitations in capturing multi-periodicity. Existing methods are either limited by one-dimensional input sequences, making it difficult to capture multi-periodicity, or require complex feature engineering [27]. This research employs the WECN framework to model tidal current time series, as it can fully utilize periodic information for forecasting. The WECN framework uses the TimesNet model proposed by Wu et al. as the infrastructure of the predictor. Wu et al. [23] observed that the time point variation for each period is influenced by the temporal patterns of its neighboring region and changes in its adjacent periods. Therefore, the two-dimensional changes in the time series can be represented as intra-period and inter-period variations.
To uniformly represent intra-period and inter-period variations, the original one-dimensional time series is transformed into a set of two-dimensional tensors based on multiple periods. The WECN framework uses visual backbone networks to capture the two-dimensional changes in the time series and achieve a unified model of intra-period and inter-period variations, as shown in Figure 3.
Figure 3: The WECN captures intra-period and inter-period variations and uses 2D convolutional kernels to process both types of variations. The time series case in the figure is selected from the used tidal current speed dataset.
First, the original input $\textbf{{X}}_{\text{1D}}$ is projected into deep features through an embedding layer, represented as:
$\textbf{{X}}_{\text{1D}}^{0}=\text{Embed}(\textbf{{X}}_{\text{1D}}),$ (4)
where $\text{Embed}(\cdot)$ represents the embedding layer. $\textbf{{X}}_{\text{1D}}^{0}$ is a two-dimensional array of shape $(T,d)$, representing the zeroth layer features in TimesNet. Here, $d$ is a hyperparameter to be optimized, representing the feature space dimension.
The adjacent TimesBlock structures are connected using the Residual structure [28]. The WECN framework is illustrated in Figure 4.
Figure 4: Overall architecture of WECN.
To obtain the features of the $l$-th layer in TimesNet, the input is $\textbf{{X}}_{\text{1D}}^{l-1}$, represented as:
$\textbf{{X}}_{\text{1D}}^{l}=\text{TimesBlock}(\textbf{{X}}_{\text{1D}}^{l-1})+\textbf{{X}}_{\text{1D}}^{l-1}.$ (5)
$\text{TimesBlock}(\cdot)$ performs three steps: ascending dimension, capturing temporal 2D-variations, and adaptive aggregation.
Ascending dimension
The input of the $l$-th layer is $\textbf{{X}}_{\text{1D}}^{l-1}$, and the calculations in Equations (1) - (3) are encapsulated in the function Period $(\cdot)$, expressed as:
$\textbf{{A}}^{l-1},\{f_{1},\cdots,f_{k}\},\{p_{1},\cdots,p_{k}\}=\text{Period}(\textbf{{X}}_{\text{1D}}^{l-1}),$ (6)
where $\textbf{{A}}^{l-1}=\{A_{f_{1}}^{l-1},\cdots,A_{f_{k}}^{l-1}\}$.
Based on the frequency set $\{f_{1},\cdots,f_{k}\}$ and the corresponding period lengths $\{p_{1},\cdots,p_{k}\}$ selected by DWT, the required number of periods is computed as $q_{i}=\left[\frac{T}{p_{i}}\right]$, and the 1D time series $\textbf{{X}}_{\text{1D}}$ is transformed into multiple 2D tensors through the following equation:
$\textbf{{X}}_{\text{2D}}^{l,i}=\text{Reshape}_{p_{i},q_{i}}(\text{Padding}(\textbf{{X}}_{\text{1D}}^{l-1})),$ (7)
where $i=1,2,\cdots,k$, $\textbf{{X}}_{\text{1D}}^{l-1}$ is a two-dimensional array of shape $(T,d)$. $\text{Padding}(\cdot)$ applies zero-padding to the sequence, extending each sequence length from $T$ to $(p_{i}\times q_{i})$. $\text{Reshape}_{p_{i},q_{i}}(\cdot)$ reshapes the sequence into a 2D tensor, resulting in $\textbf{{X}}_{\text{2D}}^{l,i}$ as a two-dimensional array of shape $(p_{i}\times q_{i},d)$.
The columns and rows of $\textbf{{X}}_{\text{2D}}^{l,i}$ represent intra-period and inter-period variations corresponding to the period length. This arrangement allows 2D kernels to handle 2D-time variations easily.
Capturing temporal 2D-variations
2D-time series features can be extracted using convolution operations. Furthermore, the model uses the Inception V1 structure from GoogLeNet [29]. For the 2D tensor, the convolution process can be expressed as:
$\hat{\mathbf{X}}^{l,i}_{\text{2D}}=\operatorname{Inception}\left(\mathbf{X}^{l,i}_{\text{2D}}\right),\ i=1,2,...,k.\\$ (8)
Adaptive aggregation
The data is first transferred from 2D to 1D space with the same period size. The Trunc()($\cdot$) function truncates the padded sequence of length $(p_{i}\times q_{i})$ to its original length $T$, expressed as:
$\hat{\mathbf{X}}^{l,i}_{\text{1D}}=\text{Trunc}\left(\text{Reshape}_{1,(p_{i}\times q_{i})}\left(\hat{\mathbf{X}}^{l,i}_{\text{2D}}\right)\right),\ i=1,2,...,k.$ (9)
Next, we need to merge the $k$ different 1D representations $\{\hat{\mathbf{X}}^{l,1}_{\text{1D}},\cdots,\hat{\mathbf{X}}^{l,k}_{\text{1D}}\}$ for the subsequent layer. The amplitude $A$ reflects the relative importance of the selected frequencies and periods, which correlates with the significance of each transformed 2D tensor [30]. Therefore, the Softmax function calculates the weights. After this, the 1D representations based on amplitude are aggregated.
[…]
$\textbf{{X}}_{\text{1D}}$ is processed through each TimesBlock with residual connections. Then, the features from the final layer are projected into the output space via a fully connected layer for multi-step forecasting. Single-step forecasting is sometimes insufficient to ensure the reliability and controllability of tidal energy systems. Therefore, this paper adopts and discusses the multi-step forecasting mechanism, providing more information for future decision-making [31]. We use a recursive forecasting strategy for multi-step forecasting, where the current output forecasting serves as the input for the next step. The multi-step forecasting process can be seen in Figure 5. The multi-step forecast results belong to $\mathbb{R}^{T\times N\times\beta}$, where $\beta$ denotes the output length. The $g$-th step forecasting value for the $t$-th sampling point in the $n$-th sequence can be represented as $\hat{x}^{n}(t+g)$, where $g=1,2,...,\beta$.
Figure 5: The multi-step forecasting process of WECN.
2.3 The process of hyperparameter optimization
We recognize the importance of selecting optimal hyperparameters. To achieve this, we use the Tree-structured Parzen Estimator (TPE). TPE is a variant of Bayesian optimization suitable for high-dimensional and discrete parameter spaces. Unlike traditional Bayesian methods that rely on Gaussian processes, TPE improves efficiency by constructing probability density functions in the hyperparameter search space using Parzen window density estimation [32]. Next, we transform the forecasting problem into an optimization process, with the design of the objective function discussed in Section 2.3.1 and the application of TPE in Section 2.3.2.
2.3.1 Objection function
Typically, multi-step time series forecasting aims to minimize the Mean Squared Error (MSE). We define the following equation as the optimization objective:
$\min_{\boldsymbol{\omega}}\ y,$ (11)
s.t. $y=\frac{1}{\beta}\sum_{g=1}^{\beta}\left(\frac{1}{NT}\sum_{t=1}^{T}\sum_{n=1}^{N}\left(\hat{x}^{n}(t+g,\boldsymbol{\omega})-x^{n}(t+g,\boldsymbol{\omega})\right)^{2}\right),$
$\boldsymbol{\omega}\in\boldsymbol{\omega}_{\text{space}},$
where $\boldsymbol{\omega}$ represents a combination of hyperparameters within the hyperparameter search space, $\hat{x}^{n}(t+g,\boldsymbol{\omega})$ represents the forecasted value, $x^{n}(t+g,\boldsymbol{\omega})$ represents the observed value. $\boldsymbol{\omega}_{\text{space}}$ denotes the hyperparameter search space, and $y$ is the value of the loss function.
2.3.2 Implementation of TPE
TPE continuously searches for the optimal solution of the objective function in a tree-structured parameter space. Here, $\boldsymbol{\omega}$ represents any hyperparameter combination within the hyperparameter search space $\boldsymbol{\omega}_{\text{space}}$, and $y$ is the loss function value under that hyperparameter combination. The optimization process of TPE is illustrated in Algorithm 1, which involves four main steps:
Initial distribution definition
Perform a random search in the initial iteration phase to help establish the initial probability distribution.
Observation splitting
Divide the hyperparameter combinations into two groups based on their performance metrics. The first group includes “good” configurations, denoted as $\chi(\boldsymbol{\omega})$, and the second group includes “bad” configurations, denoted as $\gamma(\boldsymbol{\omega})$. The performance threshold $y^{*}$ distinguishes two groups. The groups are updated after evaluating the performance of new hyperparameter combinations.
Probability modeling
Use Parzen window estimation to build probability density functions for the configurations $\chi(\boldsymbol{\omega})$ and $\gamma(\boldsymbol{\omega})$. Specifically, $p(\boldsymbol{\omega}\mid y<y^{*})$ is used for “good” configurations, and $p(\boldsymbol{\omega}\mid y\geq y^{*})$ is used for “bad” configurations.
Sampling strategy
In each iteration, select new hyperparameter combinations for evaluation based on the current probability density distributions. The selection strategy maximizes the Expected Improvement (EI), representing the likelihood that a new hyperparameter configuration will further reduce the loss, given the current best loss function value. It is calculated as follows:
$\operatorname{EI}(\boldsymbol{\omega})=\frac{\chi(\boldsymbol{\omega})}{\gamma(\boldsymbol{\omega})}.$ (12)
The hyperparameters that maximize EI are selected as the next set of search values, and the process returns to step 2 for iteration.
Output
When the stopping criteria for iteration are met, the best hyperparameter combination $\boldsymbol{\omega}_{best}$ is output.
Algorithm 1 Tree-structured Parzen Estimator
1:Initial distribution definition;
2:while the stopping criteria for iteration are not reached do
3:  Construct models $\chi(\boldsymbol{\omega})$ and $\gamma(\boldsymbol{\omega})$;
4:  Select new hyperparameters $\boldsymbol{\omega}$ to maximize $EI(\boldsymbol{\omega})=\frac{\chi(\boldsymbol{\omega})}{\gamma(\boldsymbol{\omega})}$;
5:  Evaluate performance of $\boldsymbol{\omega}$ to obtain $y$;
6:  Update configuration groups;
7:end
8:Return $\boldsymbol{\omega}_{best}$;
3 Experiment
3.1 Experimental setup
All experiments were conducted on a GeForce RTX 4090 GPU. We used a Linux operating system and the PyTorch deep learning framework under Python 3.8. To ensure the reproducibility of our experimental results, our core Python packages were aligned with the Time Series codebase, such as numpy==1.23.5 and pandas==1.5.3. The measurements are sampled at 1 Hz, and one time step equals 1 s throughout the paper. In the optimization experiments, to balance exploration of the hyperparameter space and computational efficiency, we set the stopping criterion for optimization to a maximum of 100 evaluations. Among these evaluations, the hyperparameter configuration that achieves the lowest validation loss is selected as the optimal setup for our final model.
3.2 Datasets
The Orkney Islands in the United Kingdom, located in the North Atlantic north of Scotland, are renowned for their abundant tidal energy resources and are known as an “ecological paradise” in the marine energy sector. Their unique geographical location provides real marine environmental conditions for the tidal energy test site at the European Marine Energy Centre, including various factors such as tides, ocean currents, and water depth. These data are crucial for verifying the performance of tidal energy technology in actual environments.
The tidal data used in this paper were obtained from ReDAPT project, originating from a tidal energy test site near the Fall of Warness on one of the Orkney Islands. The location can be seen in Figure 6. The TCS data were collected from four-beam Acoustic Doppler Current Profilers instruments deployed on the seabed gravity mooring frames at the tidal energy test site. To comprehensively capture the three-dimensional flow dynamics that determine the magnitude and direction of tidal currents, this study selects the velocity components in the X, Y, and Z directions as forecasting targets. Multi-level forecasting is performed at depths of 22 m, 23 m, and 24 m below sea level. The recordings span 7 days, with data collected at 1 s intervals.
Figure 6: Location of the tidal energy test site near Fall of Warness, UK.
The original data are stored using .mat file format. We first perform preprocessing. Due to malfunctions in reading instruments or interruptions caused by sudden movements near recording sensors, the data contain missing values and outliers. For consecutive missing segments, we substitute a small positive constant on the normalized scale $\varepsilon=0.01$, treating it as a near-zero measurement noise floor to maintain sequence continuity rather than interpreting them as true zeros [33]. One method to remove outliers is to pre-filter high-frequency content in the data [16]. This research uses an adaptive filter to remove high-frequency noise and erroneous data from the dataset. Table 2 shows statistics of the data at different water depths.
BSL Direction Mean Median Std
22m X -0.0633 -0.1798 1.3070
Y 0.0844 -0.0716 1.3545
Z 0.0074 0.0072 0.0699
23m X -0.0479 -0.1769 1.3272
Y 0.0852 -0.0731 1.3656
Z 0.0075 0.0072 0.0684
24m X -0.0343 -0.1754 1.3455
Y 0.0865 -0.0804 1.3771
Z 0.0077 0.0069 0.0671
Table 2: BSL stands for Below Sea Level. Mean denotes the mean values of the time series, Median denotes the intermediate value of the time series, Std is the standard deviation of the time series. The unit of tidal current velocity is meters per second (m/s).
We split the data into training, testing, and validation sets with proportions of 70%, 20%, and 10%, respectively. The validation set is designed to adjust the hyperparameters. We select TCS for a case study to provide a practical dataset and method application example.
3.3 Evaluation metrics
This subsection describes the performance metrics adopted to evaluate the forecasting capabilities of the proposed framework. Specifically, Mean Absolute Error (MAE), Mean Square Error (MSE), Mean Absolute Percentage Error (MAPE) are used as follows:
$\text{MAE}=\frac{1}{T}\sum_{t=1}^{T}|\hat{x}^{n}(t)-x^{n}(t)|$ (13)
$\text{MSE}=\frac{1}{T}\sum_{t=1}^{T}(\hat{x}^{n}(t)-x^{n}(t))^{2}$ (14)
$\text{MAPE}=\frac{1}{T}\sum_{t=1}^{T}\left|\frac{\hat{x}^{n}(t)-x^{n}(t)}{x^{n}(t)}\right|\times 100\%$ (15)
where $x^{n}(t)$ represents the observations in the datasets, $\hat{x}^{n}(t)$ is the forecasting results of the model, and $T$ denotes the number of forecasting times.
This model’s training utilizes a typical supervised learning framework. The metrics for the three directions at the same depth are averaged. The objective function of the optimization method is to minimize the average value of MSE.
Figure 7: The stacked bar chart represents the metric values of different models for 10-step forecasting. 22m indicates below sea level 22m.
3.4 Performance comparisons of different models
We conduct a comparative analysis against existing methods.
Baselines
We select the following models as baselines: XGBoost [34], LSTM [35], GRU [36], Transformer [37], Reformer [38], iTransformer [39], Autoformer [40], Pyraformer [41], DLinear [42], FEDformer [43] and FiLM [44]. The comparison models’ hyperparameter design refer to the work [45].
3.5 Performance comparisons of different models
We conduct a comparative analysis against existing methods.
Table 3: Metric values of different models for 30-step forecasting.
BSL 22m 23m 24m
Metrics MAE MSE MAPE (%) MAE MSE MAPE (%) MAE MSE MAPE (%)
XGBoost [34] 0.2546 0.1190 923.65 0.2562 0.1203 899.87 0.2573 0.1212 922.76
LSTM [35] 0.2096 0.1929 979.73 0.2398 0.2147 979.39 0.2140 0.2987 1009.88
GRU [36] 0.2877 0.1359 1452.25 0.2789 0.1283 1330.46 0.2820 0.1327 1349.83
Transformer [37] 0.1359 0.0779 102.20 0.1363 0.0779 130.76 0.1363 0.0743 150.31
Reformer [38] 0.1659 0.1031 120.16 0.1709 0.1095 147.18 0.1710 0.1077 182.59
iTransformer [39] 0.1213 0.0784 109.02 0.1206 0.0786 129.83 0.1193 0.0775 204.86
Autoformer [40] 0.2562 0.2351 257.86 0.2896 0.3117 308.84 0.2235 0.1891 264.44
Pyraformer [41] 0.1306 0.0742 96.33 0.1287 0.0724 118.99 0.1272 0.0719 178.73
DLinear [42] 0.1562 0.1061 124.45 0.1557 0.1063 132.49 0.1538 0.1050 185.70
FEDformer [43] 0.1357 0.0706 122.36 0.1335 0.0701 143.40 0.1341 0.0700 194.76
FiLM [44] 0.1536 0.1108 136.99 0.1528 0.1109 147.70 0.1512 0.1096 211.73
WECN (Ours) 0.1257 0.0701 100.53 0.1244 0.0685 121.10 0.1174 0.0679 125.66
  • 11footnotemark: 1
    BSL stands for Below Sea Level.
  • 22footnotemark: 2
    Gold represents the best performance, Silver represents the second best, Bronze represents the third best.
Table 4: Metric values of different models for 60-step forecasting.
BSL 22m 23m 24m
Metrics MAE MSE MAPE (%) MAE MSE MAPE (%) MAE MSE MAPE (%)
XGBoost [34] 0.3584 0.2222 996.64 0.3601 0.2235 1024.54 0.3612 0.2244 966.57
LSTM [35] 0.3327 0.1683 1508.65 0.3582 0.1705 1499.47 0.3503 0.1651 1558.95
GRU [36] 0.3285 0.1792 1791.93 0.3238 0.1808 1508.45 0.3039 0.1659 1620.99
Transformer [37] 0.2129 0.1656 135.84 0.2113 0.1657 176.44 0.2106 0.1645 250.67
Reformer [38] 0.2280 0.1831 156.18 0.2273 0.1840 181.79 0.2261 0.1868 246.09
iTransformer [39] 0.2026 0.1795 166.77 0.2020 0.1804 185.08 0.2000 0.1786 230.24
Autoformer [40] 0.2714 0.2641 265.11 0.2912 0.2911 280.11 0.2622 0.2543 262.56
Pyraformer [41] 0.2068 0.1618 134.66 0.2012 0.1604 156.04 0.2007 0.1596 227.37
DLinear [42] 0.2320 0.1927 165.81 0.2319 0.1937 180.28 0.2294 0.1920 229.17
FEDformer [43] 0.2113 0.1760 174.93 0.2093 0.1767 195.22 0.2072 0.1750 296.24
FiLM [44] 0.2209 0.1997 181.47 0.2205 0.2010 199.22 0.2186 0.1995 255.24
WECN (Ours) 0.2171 0.1633 134.49 0.1967 0.1534 165.74 0.1897 0.1606 229.90
  • 11footnotemark: 1
    Refer to the table 3 for the abbreviations and color codes.
Table 5: Tables of metrics for the ablation experiment of the DWT module.
Model Data MAE MSE MAPE (%)
WECN-FFT 23m 0.0323 0.0043 30.95
$23\mathrm{m}_{\mathrm{new}}$ 0.0371 $\uparrow$ 0.0115 $\uparrow$ 38.42 $\uparrow$
WECN 23m 0.0245 0.0029 25.74
$23\mathrm{m}_{\mathrm{new}}$ 0.0244 $\downarrow$ 0.0027 $\downarrow$ 25.38 $\downarrow$
  • 11footnotemark: 1
    $\uparrow$ represents an increase in forecasting error, and $\downarrow$ represents a decrease in forecasting error.
Table 6: Tables of metrics for the ablation experiment of the TPE module.
Metrics WECN w/o TPE WECN-RS WECN
MAE 0.081 0.065 0.0245
MSE 0.027 0.0047 0.0029
MAPE (%) 82.9 59.4 24.74
  • 11footnotemark: 1
    The bolded metrics indicate the best performance across different models.
  • 22footnotemark: 2
    w/o stands for without.
In this research, we design experiments to analyze the interaction between multi-periodicity and forecasting accuracy in multi-step forecasting. The evaluation first examines the model’s forecasting ability at 10-step, considering it a benchmark for short-term forecasting. Then, the analysis extends to longer multi-step forecasts, specifically the 30-step and 60-step. This progression allows us to examine whether and how the model’s performance changes with increasing time steps.
Figure 8: Radar chart of the ablation experiment for the DWT module. A represents WECN-FFT with an output length of 10. B represents WECN-FFT with an output length of 30. C represents WECN-FFT with an output length of 60. D represents WECN with an output length of 10. E represents WECN with an output length of 30. F represents WECN with an output length of 60.
Figure 7 employs a stacked bar chart to illustrate the four evaluation metrics obtained in the 10-step forecast, while Tables 3 and 4 summarize the results for the 30-step and 60-step forecasts, respectively. The reported values represent the averages computed across different directions of tidal current speeds. The WECN achieves a MAE of 0.0250, MSE of 0.0028, and MAPE of 36.23%. Compared to the best-performing model other than WECN, the MAE and MSE decrease by 0.23% and 0.37%, respectively, while MAPE increases by 7.57%. Relative to the second-best model, the reductions in MAE and MSE are 3.47% and 6.66%, respectively, accompanied by a 2.08% increase in MAPE. Overall, WECN achieves the best results in most cases. These findings demonstrate that explicitly modeling the multi-periodic characteristics of tidal currents substantially enhances forecasting accuracy. Moreover, the proposed framework maintains strong performance across datasets collected at different water depths, underscoring its robust generalization capability. We attribute this superior performance to a fundamental architectural advantage. Unlike Transformer-based models, which possess only a weak inductive bias of permutation invariance and must learn temporal ordering from scratch through positional encodings, the WECN leverages the strong inductive biases of locality and translation invariance inherent in its convolutional backbone. The Discrete Wavelet Transform (DWT) module further reinforces these biases by identifying the most salient periodic components within the data. Subsequently, the 1D-to-2D reshaping operation transforms periodicity from a learnable feature into an intrinsic structural property of the input tensor. This effectively reformulates the complex task of modeling long-range temporal dependencies into the simpler objective of capturing short-range spatial dependencies—a domain in which 2D CNNs excel.
Tables 3 and 4 show that as the forecasting time steps increase, the performance of all models significantly decreases. The previous step’s forecasting is the next step’s input in multi-step forecasting. If there is an error in each forecasting step, these errors accumulate with increasing time steps, resulting in increasingly inaccurate forecasting. Despite the forecasting horizon being extended to 30 and 60 steps, WECN effectively mitigates the degradation in predictive accuracy to a certain extent. This robustness can be attributed to WECN’s unique architectural design, which transforms the cascading temporal errors typical of 1D models into a more controllable spatial error diffusion, thereby ensuring stability in multi-step forecasting. Table 3 indicates that in the 30-step forecasting task, compared to the best model other than WECN, MAE increases by 1.74%, MSE decreases by 1.99%, and MAPE decreases by 9.12%. Compared to the second-best model other than WECN, MAE decreases by 4.89%, MSE decreases by 5.49%, and MAPE decreases by 12.20%. Table 4 shows that the model continues to perform well in the 60-step forecast, especially with data at 23m below sea level. Among the three datasets, the data at 23m exhibits the strongest non-stationarity. Multi-periodicity includes non-stationarity within the series. Therefore, the results validate the framework’s effective learning of multi-periodicity.
3.6 Ablation studies
We conduct ablation experiments on the WECN framework to determine the impact of individual components on the framework’s forecasting performance. Specifically, we examine the effects of the DWT and TPE modules.
3.6.1 Ablation study of the DWT module
First, we validate the impact of the DWT module on forecasting performance. The DWT module was ablated. We used data at 23m below sea level to compare model performance across different output steps. The WECN without the DWT module uses the Fast Fourier Transform (FFT) to learn the periodic features of the sequence, which we refer to as WECN-FFT. Since the DWT module captures the local periodicity within the sequence, replacing it increases the forecasting error metrics for WECN-FFT. The results demonstrate the advantage of the DWT module in learning multi-periodicity. Figure 8 shows the performance differences between the two periodicity-capturing methods.
To further verify the framework’s ability to capture local periodicity, we added sinusoidal local periodicity to the original data. To control locality, a Gaussian window function was used. The Gaussian window was generated based on the relationship between time, the center offset, and the window width. Then, the Gaussian window was multiplied by a sine wave to create a locally periodic signal, which was added to the original time series data, resulting in a new dataset named $23\mathrm{m}_{\mathrm{new}}$. Next, the WECN and WECN-FFT models are used to forecast both the original and new datasets. Moreover, we optimize the hyperparameters for each model separately. The performance metrics for forecasting are shown in Table 5.
The results indicate that the WECN-FFT model exhibits a decrease in prediction accuracy when dealing with data containing stronger local periodicity ($23\mathrm{m}_{\mathrm{new}}$). In contrast, the WECN framework improves prediction accuracy when handling the $23\mathrm{m}_{\mathrm{new}}$ data. The results validate the role of the DWT module. This further indicates that the high-frequency components retained in our data after being processed by the adaptive filter are meaningful and learnable patterns, not just aliased noise.
3.6.2 Ablation study of the TPE module
After analyzing the contribution of the DWT module to the framework’s predictive accuracy, we conducted the following comparative experiment to analyze further the impact of the hyperparameter optimization method on forecasting results:
  • Using Random Sampler (i.e., WECN-RS): We replaced the TPE parameter optimization method in the model with a random sampling method, which randomly selects hyperparameter values. The results showed a decline in predictive performance, indicating that the random sampling method could not effectively search for the optimal hyperparameter combination, thereby reducing the model’s generalization performance.
  • Using Default Parameter Combinations (i.e., WECN w/o TPE): We replaced the TPE parameter optimization method in the model with default parameter combinations. The results also showed a decline in forecasting performance, suggesting that the default parameter combinations could not adapt to different datasets and scenarios, limiting the model’s flexibility.
Table 6 presents the forecasting performance of the model under different hyperparameter optimization methods. The results show that the model achieved the lowest metric values using the TPE parameter optimization method. The results confirm the effectiveness of the proposed method.
4 Conclusion
In this research, we developed a multi-periodicity learning framework that is capable of learning local periodicity. We extract dominant periods via DWT and feed them into the predictor without reconstructing components. Then, the TPE algorithm is integrated to optimize hyperparameters, forming the Wavelet-Enhanced Convolutional Network (WECN) framework for tidal current speed forecasting. Numerical results demonstrate that the proposed framework attains a 10-step average MAE of 0.025. Relative to the strongest baseline, error reductions range from 0.23% to 6.66%. In practical applications, tidal current time series exhibit multi-periodicity arising from the interaction of different tidal components and local flow conditions. The proposed framework models these multi-periodic features, enabling adaptation to the complex and dynamic patterns in real marine environments, which holds significant importance for practical engineering applications. Regarding generalization ability, the framework demonstrates applicability across different datasets and can be used for similar multi-periodicity time series tasks. The results indicate that the framework can provide valuable support for integrating tidal energy into the electrical grid.
Despite its promising performance, the proposed framework still faces certain limitations that need to be addressed in future research. The current model uses tidal current time series for forecasting without considering additional features. Future work will aim to improve prediction accuracy by integrating knowledge-driven approaches.
CRediT authorship contribution statement
Tengfei Cheng: Methodology, Conceptualization, Software, Writing - original draft. Yangdi Huang: Data curation. Ling Xiao: Writing - review & editing. Yunxuan Dong: Conceptualization, Investigation, Writing - original draft, Writing - review & editing.
Funding
This work was supported in part by the Department of Science and Technology of Guangxi Zhuang Autonomous Region (grant number 2024JJB170087), in part by the Department of Human Resources and Social Security of Guangxi Zhuang Autonomous Region (grant number 202401950).
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
Data will be made available on request.